OpenStudy (anonymous):

Simplify the following expression. 2x - 2y + 5z - 2x - y + 3z x - 3y + 8z 2x - 2y + 8z 3y + 8z -3y + 8z

OpenStudy (anonymous):

i got the answer : 1y + 8z but it says im wrong.....where did i go wrong?

OpenStudy (cwrw238):

its easier if you bring like terms together

OpenStudy (anonymous):

i did and the answer i got was 1y + 8z

OpenStudy (cwrw238):

2x - 2x -2y - y + 5z + 3z

OpenStudy (anonymous):

yup that is exactly how i put them ^^^^^^^^

OpenStudy (cwrw238):

go 2x - 2x = 0 - u got that right how about - 2y - y ?

OpenStudy (anonymous):

i did 2y - y = 1y

OpenStudy (cwrw238):

no - -2y - y is not - 1y , 2y - y = 1y

OpenStudy (anonymous):

wait r u saying 2y is negative?

OpenStudy (cwrw238):

-2y is negative

OpenStudy (anonymous):

how come? what makes it negative?

OpenStudy (cwrw238):